ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 254, 2023

Adolf, D. Lower consistency bounds for mutual stationarity with divergent cofinalities and limited covering

479

Ballico, E., A. Bik, A. Oneto and E. Ventura Strength and slice rank of forms are generically equal

275

Bersudsky, M. on the image in the torus of sparse points on dilating analytic curves

341

Bik, A. See E. Ballico

Burness, T. C. On soluble subgroups of sporadic groups  313

Casals-Ruiz, M., A. Garreta, I. Kazachkov and J. de la Nuez González Simple groups with infinite verbal width and the same positive theory as free groups

39

Dong, C.-P. On the Helgason–Johnson bound  373

Erkurşun-Özcan, N. and F. Mukhamedov An extension of the zero-two law on L_{p}-spaces

461

Ferrara, M. and M. Trombetti The pro-norm of a profinite group  399
Frisch, J. and F. Shinko A dichotomy for Polish modules  97

Garreta, A. See M. Casals-Ruiz Giambruno, A., A. Ioppolo and D. La Mattina Trace codimen-

sions of algebras and their exponential growth

431

Hoffmann, D. M., C.-M. Tran and J. Ye Stable formulas in ordered structures

293

Ioppolo, A. See A. Giambruno Kazachkov, I. See M. Casals-Ruiz Kosz, D., M. Mirek, P. Plewa and B. Wróbel Some remarks on

dimension-free estimates for the discrete Hardy–Littlewood maximal functions

1

Martino, F. and C. Rizzo Differential identities and varieties of almost polynomial growth

243

La Mattina, D. See A. Giambruno Mirek, M. See D. Kosz Moretó, A. and P. H. Tiep Nonsolvable groups have a large pro-

portion of vanishing elements

229

Mukhamedov, F. See N. Erkurşun-Özcan de la Nuez González. J. See M. Casals-Ruiz Oneto, A. See E. Ballico Papadakis, S. A. and V. Petrotou The generic anisotropy of sim-

plicial 1-spheres

141

Petrogradsky, V. Scale for codimension growth of Poisson PIalgebras

201

Petrotou, V. See S. A. Papadakis Piasecki, Ł. and J. Villada Separable Lindenstrauss spaces whose

duals do not contain weak^{∗} closed convex unbounded sets having the AFPP

87

Plewa, P. See D. Kosz Rampasso, G. C. See J. V. da Silva Ricarte, G. C. See J. V. da Silva Rizzo, C. See F. Martino Shinko, F. See J. Frisch da Silva, J. V., G. C. Rampasso, G. C. Ricarte and H. A. Vivas

Free boundary regularity for a class of one-phase problems with non-homogeneous degeneracy

155

Tiep, P. H. See A. Moretó

Tomon, I. Ramsey properties of semilinear graphs  113

Tran, C.-M. See D. M. Hoffmann Trombetti, M. See M. Ferrara Ventura, E. See E. Ballico Villada, J. See Ł. Piasecki Vivas, H. A. See J. V. da Silva Wróbel, B. See D. Kosz Ye, J. See D. M. Hoffmann Zhang, J. Reflection principles, GCH and the uniformization

properties 57